OpenStudy (anonymous):

The slope of the tangent line to the curve y(cos x)+e^y=5 at the point where x=pi/2

OpenStudy (tkhunny):

Have you considered calculating dy/dx via the implicit derivative?

OpenStudy (anonymous):

how to do that,,,?

OpenStudy (anonymous):

product rule and chain rule

OpenStudy (anonymous):

\[ y\cos (x)+e^y=5\] \[y'\cos(x)-y\sin(x)+e^yy'=0\] solve for \(y'\)

OpenStudy (anonymous):

what do i do with the y when i solve for y'?

OpenStudy (agent0smith):

y(cos x)+e^y=5 plug in x=pi/2 and solve that for y. Then plug x=pi/2 and that above y value into the y' equation.